# Algebraic Connectivity Enhances Hyperedge Specificity in the Alzheimer’s Disease Continuum

### v1

@@ Abstract
Functional MRI is a neuroimaging technique aiming at analyzing the functional activity of the brain by measuring blood-oxygen-level-dependent signals throughout the brain. The derived functional features can be used for investigating brain alterations in neurological and psychiatric disorders. In this work, we employed the hypergraph structure to model high-order functional relations across brain regions, introducing algebraic connectivity ($a(\mathcal{G})$) for estimating the hyperedge weights. The hypergraph structure was derived from healthy controls to build a common topological structure across individuals. The considered cohort included subjects covering the Alzheimer’s disease (AD) continuum, that is both mild cognitive impairment and AD patients. Statistical analysis was performed using the hyperedges’ weights as features to assess the differences across the three groups. Additionally, a mediation analysis was performed to evaluate the effectiveness and reliability of the $a(\mathcal{G})$ values, representing the functional information, as the mediator between tau-PET levels, a key biomarker of AD, and cognitive scores. The proposed approach outperformed state-of-the-art methods in identifying a larger number of hyperedges statistically different across groups. Among these, two hyperedges belonging to salience ventral attention and somatomotor networks showed a partial mediation effect between the tau biomarkers and cognitive decline. These results suggested that the $a(\mathcal{G})$ can be an effective approach for extracting the hyperedge weights, including important functional information that resides in the brain areas forming the hyperedges.
Keywords Alzheimer’s Disease  $\cdot$ Functional MRI  $\cdot$ Hypergraph  $\cdot$ Mediation Analysis  $\cdot$ Weights Computation
@@ 1 Introduction
Functional magnetic resonance imaging (fMRI) is a widely used neuroimaging technique for studying changes in blood-oxygenation-level-dependent (BOLD) signals in the brain. The BOLD signals reflect the activation or deactivation of the brain areas during time, thus allowing the study of synchronous and asynchronous regions’ temporal dynamics johnson2012brain . Functional connectivity (FC) measures can be derived starting from fMRI signals, typically computed as the Pearson correlation between the time series data of each pair of brain regions and represented as symmetric connectivity matrices. FC has been extensively employed in research for investigating brain functions and identifying functional alterations possibly associated with neurodegenerative and neuropsychiatric disorders. Usually, these studies employ FC matrices in conjunction with common statistical techniques or machine/deep learning approaches to extract meaningful information alorf2022multi ; geng2018multivariate ; plitt2015functional . From FC matrices, graph structures are extracted and used to investigate the network’s patterns. The graph is a data structure that, in the neuroscience field, can represent the brain, where each node of the graph represents a brain region, and the link between nodes is expressed as a measure of the association, typically the Pearson correlation gu2025fc ; cui2022braingb . Even if the graph-based structures and corresponding metrics and learning methods can better represent the brain functions and modulations, they still represent an approximation of a complex system due to the pairwise relationships expressed by those data structures.
@@
This bottleneck can be naturally overcome by switching to hypergraphs. A hypergraph is a data structure consisting of hyperedges gathering multiple nodes. This allows capturing high-order relations across different brain regions, providing a more flexible representation with a higher number of degrees of freedom that is better suited for the problem at hand. By grouping different regions under the same hyperedge, complex relations across brain areas can be captured, relying on ad-hoc measures for hyperedge shaping. Some works have already applied functional hypergraphs in neuroscience studies, e.g., for disease classification and brain functional analysis teng2024constructing ; wang2024multiview ; liu2023deep ; gao2020hypergraph ; xiao2019multi .
@@
However, many issues remain open, leaving room for further research. The two main challenges are i) building the hypergraph backbone structure, and ii) assigning the weights to the hyperedges. Regarding the first, different algorithms and techniques have been employed for detecting regions’ interplay and creating hyperedges composed of highly related regions starting from the fMRI time series. Previous works relied on the k-nearest neighbor algorithm to define the hyperedges and corresponding brain regions teng2024constructing ; zuo2021multimodal ; zhou2006learning . Other studies adopted clustering algorithms like k-means to learn the hypergraph structure ji2022fc ; zhou2006learning . The least absolute shrinkage and selection operator (LASSO) and Elastic Net with their related extensions are two other widely used methods that regress each time series with respect to all the others, hence extracting the functional relations across regions song2023brain ; yang2023constructing ; li2020hypernetwork ; xiao2019multi ; guo2018resting . The second challenge regards weight computation. The hyperedge weights are an essential part of the data structure since meaningful information can be extracted from those values, and subsequent analysis can be conducted starting from them. A common approach for assigning weights to the hyperedges is the Gaussian similarity kernel. A distance measure is computed between pairs of node features or time series, and then the sum or mean inside the hyperedge is calculated niu2023applications ; peng2023multiview ; gao2020hypergraph ; zhang2017joint . Another common approach to compute the hyperedge weights relies on the average or sum of the correlations across the regional time series of the same hyperedge bai2025hypergraph ; song2023brain ; shao2023analysis . Other approaches derive the weights using a learning procedure, such as optimization procedures or deep learning algorithms wang2024multiview ; ji2022fc ; cai2023discovering ; xiao2019multi .
In this work, we propose a novel method to compute hyperedge weights relying on the algebraic connectivity value computed over the brain regions data composing the hyperedges as the metric to define the hyperedge weights. The rationale for using the algebraic connectivity is due to the fact that it represents a measure of how well-connected the brain regions defining the hyperedges are. This metric allows us to compress in a unique value the information provided by the different brain areas, thus creating subgraph structures (hyperedges) which could describe differences in the functional activation across groups. To evaluate this method, we analyzed fMRI data from participants along the Alzheimer’s disease (AD) spectrum, including healthy controls (HC), individuals with mild cognitive impairment (MCI), and patients with AD. We performed a group-level statistical and mediation analysis on the hyperedge weights to assess differences in FC across groups and to assess the mediation effect of the functional data between tau-PET levels, a key biomarker of AD etiopathogenesis jack2016t , and cognitive decline.
@@
In summary, the aims of this work are threefold: i) to derive hyperedge weights employing the algebraic connectivity measure in individuals across the Alzheimer’s disease continuum (HC, MCI, AD), based on a shared hypergraph structure; ii) to assess its effectiveness in differentiating diagnostic groups in relation to high-order FC patterns; and iii) to examine whether hyperedge weights derived from algebraic connectivity mediate the relationship between tau burden and cognitive decline.
@@ 2.1 Dataset
Class Subjects Age Sex (M/F) ADAS-13 HC 310310310 73.1±8.273.1$\pm$ 8.273.1 ± 8.2 135/175 8.10±4.78.10$\pm$ 4.78.10 ± 4.7 MCI 199199199 73.8±11.073.8$\pm$ 11.073.8 ± 11.0 109/90 15.0±6.615.0$\pm$ 6.615.0 ± 6.6 AD 787878 76.7±8.176.7$\pm$ 8.176.7 ± 8.1 47/31 30.1±8.630.1$\pm$ 8.630.1 ± 8.6
A total of 587 subjects were selected and stratified along the AD continuum based on their clinical status, resulting in 310 HC, 199 MCI, and 78 AD. A subset of 217 HC subjects over the 310 HC was employed only for computing the hypergraph structure, named HC-hypergraph. In contrast, the remaining 93 HC (HC-analysis) individuals were used as a separate set only for the post-hoc analysis. Table 1 shows the demographic information of the dataset used in this work.
Rs-fMRI was acquired with the following acquisition protocol: TR/TE = 3000/30 ms, FA = 90°, FOV = 220 $\times$ 220 $\times$ 163 mm3, 3.4-mm isotropic voxel size. 200 rs-fMRI volumes were acquired in almost all subjects, with minimal variations in a small subset (e.g., 197 or 195 volumes). More details on data acquisition can be found in weiner2017alzheimer .
Standard preprocessing was applied to rs-fMRI data using FSL software (https://fsl.fmrib.ox.ac.uk/fsl/), followed by nuisance regression (6 motion parameters, white matter/CSF), pass-band filtering [0.01-0.08] Hz, scrubbing, and non-linear spatial normalization to the 2-mm MNI space DeBlasi2020 . Then, the mean regional time series of the Schaefer atlas yeo2011organization were extracted, hence obtaining 100 time series with 192 time points for each subject. Finally, each time series was z-scored, resulting in a zero mean and unitary standard deviation.
@@ 2.2 Hypergraph
The brain networks analysis can be performed through graph-based data structures, such as simple graphs and hypergraphs. A simple graph $\mathcal{G}$ is a data structure that relies on pairwise relations and can be represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{w})$, where $\mathcal{V}$ represents the set of vertices (nodes), $\mathcal{E}$ is the set of edges (connections between nodes), and $\mathbf{w}$ is the vector weights where each value is associated to each edge. Simple graphs have been widely used and investigated in brain functional analyses thanks to their ability in relating pairs of regions, observing synchronous and asynchronous activities and studying brain modulations and alterations in neurodegenerative and neuropsychiatric disorders stam2024hub ; stam2014modern . Despite their potential and widespread use, they have an important limitation when dealing with complex systems, like the human brain. In detail, the simple graphs can model only pairwise relations, thus limiting the understanding of the brain functionalities and behavior. To solve this limitation, the hypergraph, which is a complex data structure, can be suitable since each hyperedge that defines the hypergraph can connect more than two nodes. A hypergraph $\mathcal{HG}$ is defined as $\mathcal{HG}=(\mathcal{V},\mathcal{E},\mathbf{W})$, where the vertex set $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ is composed of N vertices, the hyperedge set =$\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$ is composed of M hyperedges, and $\mathbf{W}$ is the diagonal matrix of the hyperedge weights, where the $m$-th element in the diagonal corresponds to weight of the $m$-th hyperedge, $\mathbf{W}=diag(\{w(e_{1}),w(e_{2}),...w(e_{M})\}$. The hypergraph can be represented by an incidence matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$, where
@@
Additionally, the incidence matrix $\mathbf{H}$ can also be non-binary, thus expressed as a weighted incidence matrix composed of real values (not only zeros and ones) in order to give a specific importance to each node in the different hyperedges. For each node, its degree represents the importance of the node in the $\mathcal{HG}$ across the hyperedges. Similarly, the hyperedge degree represents the number of nodes belonging to the hyperedge. This information in the hypergraph can be expressed in terms of diagonal matrices, the node and hyperedge degree diagonal matrices ($\mathbf{D}_{v}\in\mathbb{R}^{N\times N}$ and $\mathbf{D}_{e}\in\mathbb{R}^{M\times M}$, respectively) defined as follows:
[…]
@@ 2.2.1 Hypergraph Structure Extraction
The $\mathcal{HG}$ structure in the form of the incidence matrix $\mathbf{H}$ was derived using LASSO regression, a well-established method for hypergraph structure computation jie2016hyper ; li2019multimodal ; xiao2019multi . LASSO regression was computed in order to find relations between the regional time series, using the $n$-th time series (region $n$) as the dependent variable and the other $N$-1 time series as the independent variables. For each time series $n$ in $1\leq n\leq N$, the optimization problem is defined as
$\underset{\boldsymbol{\alpha}_{n}}{\mathrm{min}}\frac{1}{2}||\mathbf{x}_{n}-\mathbf{X}_{n}\boldsymbol{\alpha}_{n}||_{2}^{2}\ +\ \lambda||\boldsymbol{\alpha}_{n}||_{1}$ (5)
@@
where $\mathbf{x}_{n}$ is the $n$-th time series used as the response variable, $\mathbf{X}_{n}\in\mathbb{R}^{P\times N-1}$ is the time series matrix of each subject, excluding the $n$-th time series (replaced by a zero vector), $\boldsymbol{\alpha}_{n}$ corresponds to the coefficients relating the response region with the other brain areas, and $\lambda$ is the sparsity hyperparameter used to extract a sparse representation of the $\boldsymbol{\alpha}_{n}$ vector. The $\lambda$ was optimized through 5-fold cross validation and set to 0.05.
The LASSO regression coefficients were extracted for all the centroid regions ($n-th$ region), resulting in a coefficient matrix $\mathbf{A}_{s}\in\mathbb{R}^{N\times N}$ for each subject $s$. Then, the distribution of the values in each $\mathbf{A}_{s}$ was computed and subsequently thresholded, considering the fifth percentile in order to remove the smallest and non-relevant coefficients. As further step, all $\mathbf{A}_{s}$ were binarized. After that, only the hyperedges with at least three brain regions, including the centroid region, were considered for the subsequent analysis, thus obtaining an incidence matrix $\mathbf{H}_{s}\in\mathbb{R}^{N\times M}$ for each subject.
Finally, to extract a common $\mathcal{HG}$ structure for all the individuals, the majority voting value of each cell of the incidence matrices $\mathbf{H}_{s}$ was computed, resulting in $\mathbf{H}$ that represented the $\mathcal{HG}$ structure in the form of incidence matrix, as described in Eq. (1). Figure 1 shows the pipeline used to extract the hypergraph structure.

### v2

# Algebraic Connectivity Reveals Modulated High-Order Functional Networks in Alzheimer’s Disease

@@ Abstract
Functional MRI is a neuroimaging technique that analyzes the functional activity of the brain by measuring blood-oxygen-level-dependent signals throughout the brain. The derived functional features can be used for investigating brain alterations in neurological and psychiatric disorders. In this work, we employed a hypergraph to model high-order functional relations across brain regions, introducing algebraic connectivity ($a(\mathcal{G})$) for estimating the hyperedge weights. The hypergraph structure was derived from healthy controls to build a common topology across individuals. The considered cohort for subsequent analyses included subjects covering the Alzheimer’s disease (AD) continuum, encompassing both mild cognitive impairment and AD patients. Statistical analysis and three classification tasks: HC vs AD, MCI vs AD, and HC vs MCI, were performed to assess differences across the three groups and the potential of the hyperedge weights as functional features. Furthermore, a mediation analysis was performed to evaluate the reliability of the $a(\mathcal{G})$ values, representing functional information as the mediator between tau-PET levels, a key biomarker of AD, and cognitive scores. The proposed approach identified a larger number of hyperedges statistically different across groups compared to state-of-the-art methods. The $a(\mathcal{G})$ hyperedge weights also demonstrated a higher discriminative power in all three binary classifications. Finally, two hyperedges belonging to salience/ventral attention and somatomotor networks showed a partial mediation effect between the tau biomarker and cognitive decline. These results suggested that $a(\mathcal{G})$ can be an effective approach for extracting the hyperedge weights, including important functional information that resides in the brain areas forming the hyperedges.
Keywords Alzheimer’s Disease $\cdot$ Functional MRI $\cdot$ Hypergraph $\cdot$ High-Order $\cdot$ Hyperedge $\cdot$ Weights Computation
@@ 1 Introduction
magnetic resonance imaging (fMRI) is a widely used neuroimaging technique for studying changes in blood-oxygenation-level-dependent (BOLD) signals in the brain. The BOLD signals reflect the activation or deactivation of the brain areas over time, thus allowing the study of synchronous and asynchronous regions’ temporal dynamics Johnson et al. (2012). Functional connectivity (FC) measures can be derived starting from fMRI signals, typically computed as the Pearson correlation between the time series data of each pair of brain regions and represented as symmetric connectivity matrices. FC has been extensively employed in research for investigating brain functions and identifying functional alterations possibly associated with neurodegenerative and neuropsychiatric disorders. Typically, these studies employ FC matrices in conjunction with common statistical techniques or machine/deep learning approaches to extract meaningful information Alorf and Khan (2022); Geng et al. (2018); Plitt et al. (2015). From FC matrices, graph structures are extracted and used to investigate the network’s patterns. The graph is a data structure that, in the neuroscience field, can represent the brain, where each node of the graph represents a brain region, and the link between nodes is expressed as a measure of the association, typically the Pearson correlation Gu et al. (2025); Cui and others (2022). Even if graph-based structures, corresponding metrics, and learning methods can better represent brain functions and modulations, they still represent an approximation of a complex system due to the pairwise relationships expressed by these data structures.
@@
This bottleneck can be naturally overcome by switching to hypergraphs. A hypergraph is a data structure consisting of hyperedges gathering multiple nodes. This enables the capture of high-order relations across different brain regions, providing a more flexible representation with a higher number of degrees of freedom. The hyperedges reflect functionally synchronized regions over time, representing distinct local functional subnetworks that go beyond the common functional networks Gu et al. (2017). By grouping different regions under the same hyperedge, complex relations across brain areas can be captured, relying on ad-hoc measures for hyperedge shaping.
Some works have already applied functional hypergraphs in neuroscience studies, e.g., for disease classification and brain functional analysis Teng et al. (2024); Wang et al. (2024); Liu and others (2023); Gao et al. (2020); Xiao and others (2019).
@@
However, many issues remain open, leaving room for further research. The two main challenges are (i) building the hypergraph backbone structure, and (ii) assigning the weights to the hyperedges. Regarding the first, various algorithms and techniques have been employed for detecting regions’ interplay and creating hyperedges composed of highly related regions starting from the fMRI time series. Previous works relied on the k-nearest neighbor algorithm to define the hyperedges and corresponding brain regions Teng et al. (2024); Zuo et al. (2021); Zhou et al. (2006). Other studies adopted clustering algorithms like k-means to learn the hypergraph structure Ji et al. (2022); Zhou et al. (2006). The least absolute shrinkage and selection operator (LASSO) and Elastic Net, with their related extensions, are two other widely used methods that regress each time series with respect to all the others, hence extracting the functional relations across regions Song et al. (2023); Yang et al. (2023); Li and others (2020); Xiao and others (2019); Guo et al. (2018). The second challenge regards weight computation. The hyperedge weights are crucial components of the data structure, as meaningful information can be extracted from these values, and subsequent analysis can be conducted based on them. A common approach for assigning weights to the hyperedges is the Gaussian similarity kernel. A distance measure is computed between pairs of node features or time series, and then the sum or mean inside the hyperedge is calculated Niu and Du (2023); Peng et al. (2023); Gao et al. (2020); Zhang et al. (2017). Another common approach to compute the hyperedge weights relies on the average or sum of the correlations across the regional time series of the same hyperedge Bai et al. (2025); Song et al. (2023); Shao et al. (2023). Other approaches derive the weights using a learning procedure, such as optimization procedures or deep learning algorithms Wang et al. (2024); Ji et al. (2022); Cai et al. (2023); Xiao and others (2019).
In this work, we propose a novel method to compute hyperedge weights, relying on the algebraic connectivity value computed over the brain regions that comprise the hyperedges as the metric to define the hyperedge weights. One of the reasons for choosing algebraic connectivity to weight the hyperedges is that this metric reflects the degree of connectivity in a graph De Abreu (2007). Specifically, higher algebraic connectivity of a hyperedge reflects both higher synchronization and stronger network integration of its regions. On the contrary, lower values indicate poor synchronization and weaker network integration. From a biological perspective, a decrease in algebraic connectivity value can be associated with a loss of network robustness (e.g., the network becomes disconnected if a few nodes or links are removed) and reduced brain activity within the brain’s networks, deviating from the optimal configuration for dynamic processing Daianu et al. (2014); de Haan et al. (2012). Hence, by comparing the same hyperedges across groups can reveal modulations in functional connectivity between subjects. Moreover, algebraic connectivity allows a fast and effective computation through an optimization algorithm. Indeed, it does not require a large amount of resources, representing a more convenient method compared to machine/deep learning models. To evaluate this method, we analyzed fMRI data from participants along the Alzheimer’s disease (AD) spectrum, including healthy controls (HC), individuals with mild cognitive impairment (MCI), and patients with AD. We performed a group-level statistical and mediation analysis on the hyperedge weights to assess differences in FC across groups and to assess the mediation effect of the functional data between tau-PET levels, a key biomarker of AD etiopathogenesis Jack Jr and others (2016), and cognitive decline. Additionally, three binary classification tasks were performed covering the full spectrum of AD to highlight the potential of the extracted hyperedge weights as functional features.
@@
In summary, the aims of this work are threefold: (i) to derive hyperedge weights employing the algebraic connectivity measure in individuals across the AD continuum (HC, MCI, AD), based on a shared hypergraph structure; (ii) to assess its effectiveness in differentiating diagnostic groups in relation to high-order FC patterns through statistical analyses and machine learning binary classification tasks; and (iii) to examine whether hyperedge weights derived from algebraic connectivity mediate the relationship between tau burden and cognitive decline.
@@ 2.1 Dataset
HC-hypergraph HC-analysis MCI AD Subjects 217 93 199 78 Age $73.9\pm 8.7$ $70.9\pm 6.5$ $73.8\pm 11.0$ $76.7\pm 8.1$ Sex (M/F) 102/115 33/60 109/90 47/31 Education $16.9\pm 2.3$ $16.8\pm 2.3$ $16.3\pm 2.6$ $15.5\pm 2.5$ ADAS-13 $8.2\pm 4.5$ $7.8\pm 4.9$ $15.0\pm 6.6$ $30.1\pm 8.6$
A total of 587 subjects were selected from the Phase 3 of ADNI (ADNI-3) and stratified along the AD continuum based on their clinical status, resulting in 310 HC, 199 MCI, and 78 AD. A subset of 217 HC subjects over the 310 HC was employed only for computing the hypergraph structure, named HC-hypergraph, and not for the subsequent analyses. In contrast, the remaining 93 HC (HC-analysis) individuals were used as a separate set only for the post-hoc analysis along with MCI and AD groups. Hence, the last three groups were not used to construct the hypergraph backbone. Table 1 shows the demographic information of the considered groups.
Rs-fMRI was acquired with the following acquisition protocol: TR/TE = 3000/30 ms, FA = 90°, FOV = 220 $\times$ 220 $\times$ 163 mm3, 3.4-mm isotropic voxel size. 200 rs-fMRI volumes were acquired in almost all subjects, with minimal variations in a small subset (e.g., 197 or 195 volumes). More details on data acquisition can be found in Weiner and others (2017).
Standard preprocessing was applied to rs-fMRI, including removal of the first five volumes, head motion and slice-timing correction, nuisance regression (six motion parameters plus their derivatives, mean white matter (WM)/CSF signals and a linear trend component), band-pass filtering [0.01-0.08] Hz, linear plus non-linear spatial normalization to the 2-mm MNI space and smoothing with a 6-mm FWHM kernel Pini et al. (2025). Then, the mean regional time series of the Schaefer 100 atlas Yeo and others (2011) were extracted, hence obtaining 100 time series with 192 time points for each subject. Finally, each time series was z-scored, resulting in a zero mean and unitary standard deviation.
@@ 2.2 Hypergraph
Brain network analysis can be performed using graph-based data structures, such as simple graphs and hypergraphs. A simple graph $\mathcal{G}$ is a data structure that relies on pairwise relations and can be represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{w})$, where $\mathcal{V}$ represents the set of vertices (nodes), $\mathcal{E}$ is the set of edges (connections between nodes), and $\mathbf{w}$ is the vector weights where each value is associated to each edge. Simple graphs have been widely used and investigated in brain functional analyses thanks to their ability in relating pairs of regions, observing synchronous and asynchronous activities, and studying brain modulations and alterations in neurodegenerative and neuropsychiatric disorders Stam (2024, 2014). Despite their potential and widespread use, they have an important limitation when dealing with complex systems, like the human brain. In detail, simple graphs can model only pairwise relations, thus limiting the understanding of brain functionalities and behavior. To solve this limitation, the hypergraph, which is a complex data structure, can be suitable since each hyperedge that defines the hypergraph can connect more than two nodes. A hypergraph $\mathcal{HG}$ is defined as $\mathcal{HG}=(\mathcal{V},\mathcal{E},\mathbf{W})$, where the vertex set $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ is composed of N vertices, the hyperedge set $\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$ is composed of M hyperedges, and $\mathbf{W}$ is the diagonal matrix of the hyperedge weights, where the $m$-th element in the diagonal corresponds to weight of the $m$-th hyperedge, $\mathbf{W}=diag(\{w(e_{1}),w(e_{2}),...w(e_{M})\}$. The hypergraph can be represented by an incidence matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$, where
@@
Additionally, the incidence matrix $\mathbf{H}$ can also be non-binary, expressed as a weighted incidence matrix composed of real values (not only zeros and ones) in order to assign a specific importance to each node in the different hyperedges. For each node, its degree represents the importance of the node in the $\mathcal{HG}$ across the hyperedges. Similarly, the hyperedge degree represents the number of nodes belonging to the hyperedge. This information in the hypergraph can be expressed in terms of diagonal matrices and, building on this, the node and hyperedge degree diagonal matrices ($\mathbf{D}_{v}\in\mathbb{R}^{N\times N}$ and $\mathbf{D}_{e}\in\mathbb{R}^{M\times M}$, respectively) can be defined as follows:
$d(v_{n})$ $=\sum_{e_{m}\in\mathcal{E}}w(e_{m})\mathrm{H}_{n,m}$ $\mathrm{for}\ 1\leq n\leq N$ (2)
$\delta(e_{m})$ $=\sum_{v_{n}\in\mathcal{V}}\mathrm{H}_{n,m}$ $\mathrm{for}\ 1\leq m\leq M$ (3)
@@ 2.2.1 Hypergraph Structure Extraction
The $\mathcal{HG}$ structure in the form of the incidence matrix $\mathbf{H}$ was derived using LASSO regression, a well-established method for hypergraph structure computation Jie et al. (2016); Li and others (2019); Xiao and others (2019). LASSO regression was computed in order to identify relationships between regional time series, considering the $n$-th time series (region $n$) as dependent variable and the other NN-1 time series as independent variables. Therefore, for each time series $n$ in $1\leq n\leq N$, the optimization problem is defined as
[…]
@@
where the constraint prevents the negative coefficients, $\mathbf{x}_{n}$ is the $n$-th time series used as the response variable, $\mathbf{X}_{n}\in\mathbb{R}^{P\times N-1}$ is the time series matrix of each subject, excluding the $n$-th time series (replaced by a zero vector), $\boldsymbol{\alpha}_{n}$ corresponds to the coefficients relating the response region with the other brain areas, and $\lambda$ is the sparsity hyperparameter used to extract a sparse representation of the $\boldsymbol{\alpha}_{n}$ vector. The $\lambda$ was optimized through 5-fold cross validation, testing different values in the range of [0.01, 0.30], with a step of 0.01. Specifically, for every $\lambda$, for every subject, and region used as centroid, the 5-fold cross validation was performed on the timepoints of the fMRI time series. This allowed to find the LASSO model that best reconstructed the centroid using the other regions as features. At the end of this procedure, $\lambda$ = 0.05 was selected based on the overall best performance, where the average mean squared error across all subjects and regions was $0.115\pm 0.02$ for training and $0.257\pm 0.05$ for validation.
The LASSO regression coefficients were extracted for all the centroid regions ($n$-th region), resulting in a coefficient matrix $\mathbf{A}_{s}\in\mathbb{R}^{N\times N}$ for each subject $s$. Then, the distribution of the values in each $\mathbf{A}_{s}$ was computed and subsequently thresholded, considering the fifth percentile to remove the smallest and non-relevant coefficients. As a further step, all $\mathbf{A}_{s}$ were binarized. After that, only the hyperedges with at least three brain regions, including the centroid region, were considered for the subsequent analysis, thus obtaining an incidence matrix $\mathbf{H}_{s}\in\mathbb{R}^{N\times M}$ for each subject. The threshold of a minimum of three regions was selected to consider all the possible high-order interactions while avoiding the pairwise connections captured by conventional graph analyses. All the extracted hyperedges were considered undirected since the regions that formed them were directly linked with the centroid region, and the same centroid region was part of the hyperedge.
Finally, to extract a common $\mathcal{HG}$ structure for all the individuals, the majority voting values across all the incidence matrices $\mathbf{H}_{s}$ were computed. In detail, for each cell, the value of 1 or 0 was selected if at least 50% of the HC-hypergraph individuals had a value of 1 or 0, respectively. The result consisted of a matrix $\mathbf{H}$ that represented the $\mathcal{HG}$ structure in the form of incidence matrix, as described in Eq. (1).
Figure 1 shows the pipeline used to extract the hypergraph structure, while Fig. 2 shows the incidence matrix $\mathbf{H}$ of the hypergraph. The resulting hypergraph $\mathcal{HG}$ is composed of 96 hyperedges with a maximum of seven regions per hyperedge, a minimum of three, and an average of 4.66 regions. The cardinality of each hyperedge was calculated using Eq. (3), which extracts the hyperedge degree. The extracted $\mathcal{HG}$ structure remained stable even if shorter time series were employed, by truncating the final portion of each fMRI signal. This stability was assessed using the Jaccard index and Pearson’s correlation between the $\mathcal{HG}$ structure extracted from the full-length time series and the shorter signals.
@@
Figure 2: Incidence matrix H of the hypergraph. H is composed of 96 hyperedges, with a maximum of 7 regions per hyperedge, a minimum of 3, and an average of 4.66 regions.